# Proxy-small objects present compactly generated categories

### v1

@@ 7. A perspective on complete intersections
Let $Q$ be a finite quiver, $k$ a field, and $I$ an ideal in $kQ$ generated by paths. Assume that the monomial algebra $\Lambda=kQ/I$ is Iwanaga-Gorenstein. According to [4, Theorem 6.3] there is an explicit map $\chi\colon\Sigma^{-p}\Lambda\longrightarrow\Lambda$ in $\mathsf{D}(\Lambda^{\mathrm{op}}\otimes_{k}\Lambda)$ such that, for all finite dimensional $\Lambda$-modules $M$ and $N$, $\operatorname{Ext}^{*}_{\Lambda}(M,N)=\mathrm{HH}^{*}(\Lambda,\operatorname{Hom}_{k}(M,N))$ is finitely generated over the subring $k[\chi]\subseteq\mathrm{HH}^{*}(\Lambda)$; that is, $\Lambda$ satisfies the so-called ${\bf Fg}$ condition. Varying $M$ and $N$ over the simple $\Lambda$-modules we obtain all simple $\Lambda$-bimodules as $\operatorname{Hom}_{k}(M,N)$. Therefore, for each simple bimodule $s$, applying the functor $\operatorname{Ext}_{\Lambda^{\mathrm{op}}\otimes_{k}\Lambda}(-,s)$ to the triangle
@@
The extension algebra $\operatorname{Ext}^{*}_{\Lambda}(k,k)$ of $\Lambda$ is finite over a subalgebra $k\langle u,v\rangle/(uv-q^{mn}vu)$ [1, 5.3]. Moreover one can compute using [3, 3.1] that, under the map
@@
Rationally elliptic spaces are “spherically complete intersection” in the sense of [10]—they can be constructed, up to rational homotopy equivalence, by starting with a point and using finitely many spherical fibrations. Therefore by [10, Theorem 10.1] they are “endomorphism complete intersection”, that is, there is a sequence of triangles of dg $C^{*}(X\times X;\mathbb{Q})$-modules
@@
such that $M_{c}$ is a perfect dg $C^{*}(X\otimes X;\mathbb{Q})$-module. Each $M_{i}$ has finite dimensional total homology, and by [10, Remark 10.2] each $g_{i}$ has strictly positive cohomological degree. Therefore, as in Example 7.8, we can recover each $M_{i-1}$ from $M_{i}$ by taking a countable homotopy colimit. It follows that $C^{*}(X)$ is proxy-small as a dg module over $C^{*}(X\times X)\simeq C^{*}(X)\otimes C^{*}(X)$ with proxy $M_{c}$, as claimed.
@@
Nevertheless, $P$ is not proxy-small. Indeed, $\operatorname{thick}(P)$ consists of the bounded complexes of injectives with finite dimensional cohomology and so $\operatorname{thick}(P)\cap\mathsf{D}^{\mathrm{perf}}(R)$ consists of complexes with both finite projective and injective dimension. The only such complex can be 00, because $R$ is not Gorenstein; see for example [7, 2.10].

### v2

@@ 7. A perspective on complete intersections
Let $Q$ be a finite quiver, $k$ a field, and $I$ an ideal in $kQ$ generated by paths. Assume that the monomial algebra $\Lambda=kQ/I$ is Iwanaga-Gorenstein. According to [5, Theorem 6.3] there is an explicit map $\chi\colon\Sigma^{-p}\Lambda\longrightarrow\Lambda$ in $\mathsf{D}(\Lambda^{\mathrm{op}}\otimes_{k}\Lambda)$ such that, for all finite dimensional $\Lambda$-modules $M$ and $N$, $\operatorname{Ext}^{*}_{\Lambda}(M,N)=\mathrm{HH}^{*}(\Lambda,\operatorname{Hom}_{k}(M,N))$ is finitely generated over the subring $k[\chi]\subseteq\mathrm{HH}^{*}(\Lambda)$; that is, $\Lambda$ satisfies the so-called ${\bf Fg}$ condition. Varying $M$ and $N$ over the simple $\Lambda$-modules we obtain all simple $\Lambda$-bimodules as $\operatorname{Hom}_{k}(M,N)$. Therefore, for each simple bimodule $s$, applying the functor $\operatorname{Ext}_{\Lambda^{\mathrm{op}}\otimes_{k}\Lambda}(-,s)$ to the triangle
@@
The extension algebra $\operatorname{Ext}^{*}_{\Lambda}(k,k)$ of $\Lambda$ is finite over a subalgebra $k\langle u,v\rangle/(uv-q^{mn}vu)$[1, 5.3]. Moreover one can compute using [4, 3.1] that, under the map
@@
Rationally elliptic spaces are “spherically complete intersection” in the sense of [11]—they can be constructed, up to rational homotopy equivalence, by starting with a point and using finitely many spherical fibrations. Therefore by [11, Theorem 10.1] they are “endomorphism complete intersection”, that is, there is a sequence of triangles of dg $C^{*}(X\times X;\mathbb{Q})$-modules
@@
such that $M_{c}$ is a perfect dg $C^{*}(X\otimes X;\mathbb{Q})$-module. Each $M_{i}$ has finite dimensional total homology, and by [11, Remark 10.2] each $g_{i}$ has strictly positive cohomological degree. Therefore, as in Example 7.8, we can recover each $M_{i-1}$ from $M_{i}$ by taking a countable homotopy colimit. It follows that $C^{*}(X)$ is proxy-small as a dg module over $C^{*}(X\times X)\simeq C^{*}(X)\otimes C^{*}(X)$ with proxy $M_{c}$, as claimed.
@@
Nevertheless, $P$ is not proxy-small. Indeed, $\operatorname{thick}(P)$ consists of the bounded complexes of injectives with finite dimensional cohomology and so $\operatorname{thick}(P)\cap\mathsf{D}^{\mathrm{perf}}(R)$ consists of complexes with both finite projective and injective dimension. The only such complex can be 00, because $R$ is not Gorenstein; see for example [8, 2.10].